OpenStudy (anonymous):

Find all numbers c that satisfy the conclusion of Rolle's Theorem for the following function. If there are multiple values, separate them with commas; enter N if there are no such values. f(x)=x^2-8x +7, [0,8]

OpenStudy (anonymous):

f(c)'=f(8)-F(0)/8-0 2c-8=7-7/8 2c-8=0 2c=8 c=4